OpenStudy (anonymous):

algebra 1 math! write the equation of the line that is purpendicular to the graph of y=1/2 x + 6 and whose y-intercept is (0, -2) write steps please (if you can)

OpenStudy (saifoo.khan):

Again? :S

OpenStudy (anonymous):

LOL sorry

OpenStudy (anonymous):

you don't have to answer

OpenStudy (saifoo.khan):

i will get.

OpenStudy (saifoo.khan):

lol.

OpenStudy (saifoo.khan):

Do u want the ans?

OpenStudy (anonymous):

yeah

OpenStudy (saifoo.khan):

This guy is helping! ;)

OpenStudy (anonymous):

y=mx+b b= y value of y intercept=-2 for perpendicular lines the slopes has to be opposite of the reciprocal so if 1 slope is a the other has to b -1/a for this 1 the slope is 1/2 so the new line slope must be: m=-2 so y=-2x-2

OpenStudy (anonymous):

The formula of a line is : \[y - y _{0} = m(x-x _{0})\] Where y & x are variables; \[y _{0} & x _{0}\] are the coordinates of a point on the straight line - (0,2) in this case and m is the slope. For a line to be perpendicular to another, their slopes must obey the relation mm'=-1. The slope of a line is the coefficient of x when the coefficient of y is 1. If the coefficient of y is not 1, divide the equation by that (coefficient of y) to make it one. Here the slope of the given equation is 1/2. put it in the relation between the slopes (mm'=-1): (1/2)m'=-1 => m' = -2 now put the given point and the slope in the main formula for the line. that gives: y-(-2)=-2x-(-2x0) => y= -2x -2